F(x)={3,x<-2
{x+2,-2<=x<2
{-2x, x>=2
F(-2)

Answer by stanbon(75887)

F(x)={3,x<-2
{x+2,-2<=x<2
{-2x, x>=2
=====================
F(-2) = -2+2 = 0
=====================
Cheers,
Stan H.
=====================